Consider the basic food items consumed by a rural household:


FOOD ITEM CONSUMPTION PRICE($)

JUNE 2021 JULY 2021 JUNE JULY

Bread 25 26 10.30 10.70

(loaves)

Cooking oil 5 6 45.00 59.40

(litres)

Rice(kg) 10 8 34.50 36.20


Meat(kg) 15 16 88.90 92.20


A). Which food item has had the lowest relative quantity of consumption increase between the months of June and July?

B). Which food item has had the lowest relative price increase between the months June and July 2021?

C). Calculate and interpret a Paasche quantity index for July 2021. 


1
Expert's answer
2021-09-06T12:45:25-0400

Given information:



a)

Form the excel table for the relative increase of Consumption:

"Relative \space increase=\frac{New \space consumption−Old \space consumption}{Old \space consumption}"



Among the food items, the item that had the lowest relative quantity of consumption increase between the months of June and July is Rice whose relative quantity of consumption increase is - 0.2(20%). Rice is the only food item whose consumption has decreased in July as compared to June.


b)

Form the excel table for the relative increase of price:

"Relative \space increase=\frac{New \space price−Old \space price}{Old \space price}"



Among the food items, the item that had the lowest relative price increase between the months of June and July is Meat whose relative price increase is 0.037(3.71%).

c)

"Paasche \space quantity \space index=\frac{∑Observation \space quantity \times Observation \space price}{∑Base \space quantity×Observation \space price} ×100 \% \\

= \frac{26×10.70+6×59.40+8×36.20+16×92.20}{25×10.70+5×59.40+10×36.20+15×92.20}×100 \%\\

=103.89 \%"

The Paasche quantity index for July 2021 is 103.89%.

Since the Paasche quantity index for July 2021 is 103.89% which is greater than 100, hence it signifies the Inflation impact.